### v1

@@ 3.3.2 Comparison with and without reward normalization in CRL and non-CRL
Reached target successfully
44.80%
10.35%
7.15%
2.50%
Terminated by collision or driving outside road
13.45%
8.80%
9.60%
6.20%
Driven successfully, but not reached the target within maximum steps
41.75%
80.85%
83.25%
91.30%
Average distance travelled
1269.80 m
837.52 m
739.12 m
603.66 m
Average executed steps
265.91
431.43
438.64
463.48
Average speed
11.38 m/s
3.80 m/s
2.99 m/s
2.14 m/s
Average energy cost
1.23 euros
0.38 euros
0.27 euros
0.20 euros
Average driver cost
3.7 euros
6 euros
6.1 euros
6.44 euros
Average tcop
4.93 euros
6.38 euros
6.37 euros
6.64 euros
@@
Reached target successfully
57.45%
49.15%
73.50%
73.25%
Terminated by collision or driving outside road
8.25%
4.45%
2.90%
3.50%
Driven successfully, but not reached the target within maximum steps
34.30%
46.40%
23.60%
23.25%
Average distance travelled
1569.99 m
1382.20 m
1794.74 m
1770.03 m
Average executed steps
252.24
294.00
210.22
206.94
Average speed
12.27 m/s
10.38 m/s
14.18 m/s
14.42 m/s
Average energy cost
1.33 euros
1.08 euros
1.68 euros
1.62 euros
Average driver cost
3.50 euros
4.08 euros
2.92 euros
2.88 euros
Average tcop
4.83 euros
5.17 euros
4.60 euros
4.50 euros
@@
Reached target successfully
7.15%
73.50%
2.50%
73.25%
Terminated by collision or driving outside road
9.60%
2.90%
6.20%
3.50%
Driven successfully, but not reached the target within maximum steps
83.25%
23.60%
91.30%
23.25%
Average distance travelled
739.12 m
1794.74 m
603.66 m
1770.03 m
Average executed steps
438.64
210.22
463.48
206.94
Average speed
2.99 m/s
14.18 m/s
2.14 m/s
14.42 m/s
Average energy cost per meter
0.00036 euros
0.00094 euros
0.00033 euros
0.00092 euros
Average driver cost per meter
0.0082 euros
0.0016 euros
0.0107 euros
0.0016 euros
Average tcop per meter
0.0086 euros
0.0026 euros
0.011 euros
0.0025 euros
@@ Acknowledgment
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. We would like to thank Nikolce Murgovski, Erik Börve and Stefan Börjesson for valuable discussions.

### v2

@@ 3.3.2 Comparison with and without reward normalization in CRL and non-CRL
Reached target successfully 44.80% 10.35% 7.15% 2.50%
Terminated by collision or driving outside road 13.45% 8.80% 9.60% 6.20%
Driven successfully, but not reached the target within maximum steps 41.75% 80.85% 83.25% 91.30%
Average distance travelled 1269.80 m 837.52 m 739.12 m 603.66 m
Average executed steps 265.91 431.43 438.64 463.48
Average speed 11.38 m/s 3.80 m/s 2.99 m/s 2.14 m/s
Average energy cost 1.23 euros 0.38 euros 0.27 euros 0.20 euros
Average driver cost 3.7 euros 6 euros 6.1 euros 6.44 euros
Average tcop 4.93 euros 6.38 euros 6.37 euros 6.64 euros
@@
Reached target successfully 57.45% 49.15% 73.50% 73.25%
Terminated by collision or driving outside road 8.25% 4.45% 2.90% 3.50%
Driven successfully, but not reached the target within maximum steps 34.30% 46.40% 23.60% 23.25%
Average distance travelled 1569.99 m 1382.20 m 1794.74 m 1770.03 m
Average executed steps 252.24 294.00 210.22 206.94
Average speed 12.27 m/s 10.38 m/s 14.18 m/s 14.42 m/s
Average energy cost 1.33 euros 1.08 euros 1.68 euros 1.62 euros
Average driver cost 3.50 euros 4.08 euros 2.92 euros 2.88 euros
Average tcop 4.83 euros 5.17 euros 4.60 euros 4.50 euros
@@
Reached target successfully 7.15% 73.50% 2.50% 73.25%
Terminated by collision or driving outside road 9.60% 2.90% 6.20% 3.50%
Driven successfully, but not reached the target within maximum steps 83.25% 23.60% 91.30% 23.25%
Average distance travelled 739.12 m 1794.74 m 603.66 m 1770.03 m
Average executed steps 438.64 210.22 463.48 206.94
Average speed 2.99 m/s 14.18 m/s 2.14 m/s 14.42 m/s
Average energy cost per meter 0.00036 euros 0.00094 euros 0.00033 euros 0.00092 euros
Average driver cost per meter 0.0082 euros 0.0016 euros 0.0107 euros 0.0016 euros
Average tcop per meter 0.0086 euros 0.0026 euros 0.011 euros 0.0025 euros
@@ 4 Conclusion
An interesting future direction would be to explore transfer learning methods for generalizing tactical decision-making to diverse traffic scenarios such as uphill, downhill, merging traffic etc. In this context, we believe that the pre-trained models from simple highway environments can serve as a strong foundation while adapting to new environments with different dynamics and constraints.
@@ Acknowledgment
This work was partially supported by Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by Knut and Alice Wallenberg Foundation. We would like to thank Nikolce Murgovski, Erik Börve and Stefan Börjesson for valuable discussions.